WELLNESS WORKSHEET 4-continued Maintenance Rate (or how long to sip a drink) Remember that the effects of alcohol will be greater when your BAC Ia tising than when yon kong alow it to fall. BAC is darectly proportional to the rate of edaryl alcobol iatake. Assuming a geeril mn nance rate (rate at which the body rids itself of alcobol) of 0.1 oz of pure alcobol per hour per 50 g body weighl you can calculate the approximale leagth of time itksouolis en okb applying the following formola 2.5 x proof of drink x volume (size in ox) of drink body weigh - time in hours per drink For example, to calculate how long it will take to metabolize one can (12 oa) of 10-grool beer for a pernce weighing 150 pounds: 2.5 x 10x12-2 hours So, in takes this 150-pound individual 2 hours to complesely metaboline one 12 oz cam of 10-proof beer your favorite three drinks (or choose three of the examples from the peevicus page), ad une tdias formula to calculate your maintenance rate for each drink 2. x)x 3.xx hours/drink sober up, the only remedy that works is to stop drinking and allow time. For any given type of drink the amount of time would be the number of drinks you have consumed maliplid by your maintenance rate for that drink. For the example given above, if the 150-pound individal had consumed three 12 oz cans of 10-proof beer, he or she would have to wait 6 hours before the akcobol would be metabolized. Cadeoulate bhe amount of time that would have to elapse for you to metabolize all the alcobal if you bad cosumed three o one of the types of drinks you calculated a maintenance rate for above In Case of Excess 3x( )= hours level, your answer here indicates the mumber of bours you sbouald wait before driving Given this
Answer #1

* Examples of 3 drinks which are used to calculate the maintainance dose

   1. Volume of 14 ox of 10 proof beer for a person with a body weight of 160 pounds

                       Formula = 2.5 * proof of drink * volume of drink (size in ox) / Body weight.

                                       = 2.5 * 10* 14 / 140

                                       = 2.5 hours/ drink

    2. Volume of 10 ox of 20 proof beer for a person with a body weight of 140 pounds.

                           As per above formula,

                                     = 2.5* 20 * 10 /140

                                     = 3.57 hours/ drink

   3. Volume of 20 ox of 10 proof beer for a person with body weight of 150 pounds.

                          As per above formula

                                     = 2.5 * 10 * 20 / 150

                                    = 3.33 hours/ drink

* In case of excess

           * If the individual with a body weight of 160 pounds consumed volume of three 14 ox cases

             of 10 proof beer then

                        Maintainance rate of the drink is,

                                Formula = Number of drinks consumed * Maintainance rate for that drink.

                                                          = 3 * 2.5 hours/ drink

                                                          = 7.5 hours / drink